# A Systematic Search for AGN Flares in ZTF Data Release 23

### v2

@@ Abstract
software: celerite (foreman-mackeyFastScalableGaussian2017), astropy (astropycollaborationAstropyCommunityPython2013; astropycollaborationAstropyProjectBuilding2018; astropycollaborationAstropyProjectSustaining2022), matplotlib (hunterMatplotlib2DGraphics2007), numpy (harrisArrayProgrammingNumPy2020), seaborn (Waskom2021), pandas (mckinney-proc-scipy-2010)
@@ 1 Introduction
Active galactic nuclei (AGNs) are extremely luminous sources powered by accretion onto supermassive black holes (SMBHs), and they exhibit significant variability across different wavelengths (ulrichVARIABILITYACTIVEGALACTIC1997; kellyAREVARIATIONSQUASAR2009). The optical light curves of AGNs often show stochastic brightness fluctuations on timescales from days to years, commonly modeled by a one-dimensional damped random walk (macleodMODELINGTIMEVARIABILITY2010; kozlowskiRevisitingStochasticVariability2016). These fluctuations can be attributed to various processes occurring near the supermassive black hole, such as accretion disk instabilities (treveseQuasarSpectralSlope2002), disk inhomogeneities (dexterQUASARACCRETIONDISKS2010; caiSimulatingTimescaleDependentColor2016; caiEUCLIAExploringUV2018), or X-ray reprocessing (kubotaPhysicalModelBroadband2018), but the primary origin is still under debate. In addition to the general optical variability of AGN over different timescales, a subset of these systems exhibit transient and extreme events, known as AGN flares (grahamUnderstandingExtremeQuasar2017; lawrenceSlowblueNuclearHypervariables2016). These events represent a significant departure from the typical stochastic variability of AGN, yet their origins remain largely uncertain.
@@
A growing body of theoretical work suggests that many of these extreme flaring events may originate from stellar and compact objects embedded in AGN accretion disks, broadly referred to as accretion-modified stars (AMSs, wangAccretionmodifiedStarsAccretion2021a; wangAccretionmodifiedStarsAccretion2021; liuAccretionmodifiedStarsAccretion2024). Within such dense and gas-rich environments, stars may evolve rapidly and end their lives as supernovae (liCorecollapseSupernovaExplosions2023), or be tidally disrupted by the central SMBH (ryuTidalDisruptionsMainsequence2020) or stellar-mass black holes (yangTidalDisruptionStellarmass2022). These processes contribute to the metal enrichment of the AGN disk (wangStarFormationSelfgravitating2023) and the formation of compact objects. A large population of compact objects can form binaries (rowanBlackHoleBinary2023; wangStellarBHPopulation2024) and undergo mergers (mckernanBlackHoleNeutron2020). The resulting coalescences can produce gravitational waves and potentially be accompanied by observable electromagnetic signals, including transients from white dwarf collisions (zhangElectromagneticSignaturesWhite2023), kilonova emission from neutron star-neutron star or neutron star-black hole mergers (renInteractingKilonovaeLonglasting2022), and flares from binary black hole mergers (grahamCandidateElectromagneticCounterpart2020; grahamLightDarkSearching2023)
@@
These scenarios offer a unified framework for interpreting diverse AGN flaring phenomena and highlight the importance of searching for AGN flares as a means to probe the AMS populations. In addition, AGN flares serve as valuable diagnostics of accretion physics. Their timescales and amplitudes may constrain the properties of AGN disks and act as important probes for accretion rates (ruanUNDERSTANDINGCHANGINGLOOKQUASARS2016; grahamUnderstandingExtremeQuasar2017). Multi-epoch spectroscopy during the period of the flares further enables reverberation mapping to get the location of the emission region, helping constrain the structure of the AGN disk (payneChandraHSTSTIS2023; cabreraSearchingElectromagneticEmission2024a).
@@
Recognizing their scientific potential, several systematic searches have been conducted to identify AGN flares (lawrenceSlowblueNuclearHypervariables2016; grahamUnderstandingExtremeQuasar2017), utilizing data from large surveys like Panoramic Survey Telescope and Rapid Response System (Pan-STARRS, kaiserPanSTARRSWidefieldOptical2010), Sloan Digital Sky Survey (SDSS, sesarExploringVariableSky2007), and Catalina Real-time Transient Survey (CRTS, drakeFirstResultsCatalina2009). These studies have primarily focused on detecting long-duration, high-amplitude variability events, primarily due to the limited observational cadence of these surveys.
@@
The emergence of high-cadence time-domain facilities, such as the Zwicky Transient Facility (ZTF, bellmZwickyTransientFacility2019; grahamZwickyTransientFacility2019), the Wide Field Survey Telescope (WFST, wangScience25meterWide2023), and the Vera C. Rubin observatory (collaborationScienceDrivenOptimizationLSST2017; ivezicLSSTScienceDrivers2019), would be able to capture much more AGN flares across a wide range of timescales. These advances would facilitate the discovery and characterization of AGN flares, offering deeper insights into the physical mechanisms driving such outbursts.
@@
To fully utilize the data from modern high-cadence surveys, recent studies have proposed the use of Gaussian Processes (GPs) for the detection of AGN flares (mclaughlinUsingGaussianProcesses2024). grahamLightDarkSearching2023 applied a GP-based technique to separate flaring events from typical AGN variability, while heTracingLightIdentification2025 further demonstrated that applying such methods to larger datasets significantly improves the reliability of flare identification. Motivated by these developments, we carry out a systematic search for AGN flares in ZTF Data Release 23 (DR23) using a GP-based approach, aiming to construct the most comprehensive catalog of such events to date.
@@
The outline of this paper is as follows. In Section 2, we describe the ZTF dataset and the AGN catalogs used in this work. In Section 3, we introduce the flare detection method, including Gaussian Process modeling and the selection criteria used to construct two catalogs: the AGN Flare Coarse Catalog (AGNFCC) and the AGN Flare Refined Catalog (AGNFRC). In Section 4, we present the statistical properties of the detected flares. In Section 5, we discuss their potential origins. Finally, we summarize our main findings in Section 6.
2 data
2.1 ZTF DR23
@@
The Zwicky Transient Facility (ZTF) is a wide-field time-domain survey utilizing a 47 deg2 field-of-view camera mounted on the Palomar 48-inch Samuel Oschin Schmidt telescope. Designed to systematically explore transient and variable phenomena with high cadence and sensitivity (bellmZwickyTransientFacility2018; grahamZwickyTransientFacility2019), ZTF has been operating since March 2018, covering approximately $3\pi$ sr across the Northern sky (bellmZwickyTransientFacility2019). For the Northern sky of declination $31^{\circ}$, ZTF operates two primary observing programs: the Survey, which targets fields with $|b|<7^{\circ}$ at a 1-day cadence, and the Northern Sky Survey, which observes fields with $|b|>7^{\circ}$ at a 3-day cadence. Both surveys acquire data in the g and r bands, providing dual-band light curves for each source.
@@ 2.2 AGN Catalogs
Million Quasars Catalog (Milliquas, v8, fleschMillionQuasarsMilliquas2023): A comprehensive catalog of spectroscopically confirmed AGNs and high-confidence AGN candidates ($p_{\mathrm{QSO}}\geq 99\%$), compiled from all published papers up to 2023 June 30. It contains 1,021,800 sources, with $90\%$ having spectroscopic redshifts and the remainder utilizing photometric redshifts. This catalog incorporates astrometric corrections to historical positional errors and integrates high-precision multi-survey data, delivering the most precise locations.
DESI AGN Catalog (collaborationDataRelease12025): The Dark Energy Spectroscopic Instrument (DESI) has conducted the most extensive extragalactic spectroscopic redshift survey to date. Its first Data Release (DR1) provides high-confidence redshifts for approximately 18 million objects across over 9,000 deg2. In this work, we use the subset of objects labeled with SPECTYPE = ‘QSO’ in the DESI value-added catalog [^1] https://data.desi.lbl.gov/public/dr1/vac/dr1/agnqso (Juneau et al 2025, in preparation), which includes a total of 1,772,694 AGNs. This represents the largest spectroscopic AGN sample available to date, making it a valuable source for statistical studies of AGN.
LAMOST AGN Catalog (aiLargeSkyArea2016; dongLargeSkyArea2018; yaoLargeSkyArea2019; jinLargeSkyArea2023): A set of spectroscopically confirmed AGNs from the Large Sky Area Multi-Object Fiber Spectroscopic Telescope (LAMOST) quasar survey conducted between 2011 and 2021. The full sample contains 55,636 quasars, of which over 30,000 are independently discovered. All spectra are recalibrated using SDSS or Pan-STARRS photometry, and the key spectral properties, including emission line widths, continuum luminosities, and black hole masses, are homogeneously measured.
Quaia AGN Catalog (storey-fisherQuaiaGaiaunWISEQuasar2024): An all-sky catalog constructed from sources identified as quasar candidates in Gaia Data Release 3 (bailer-jonesGaiaDataRelease2023), initially comprising 6,649,162 candidates with redshift estimates derived from the low-resolution blue photometer/red photometer spectra. To improve the purity of the sample, these candidates are cross-matched with the unWISE infrared catalog (schlaflyUnWISECatalogTwo2019) and refined by applying cuts based on color and proper motion to reduce contamination. Redshift estimates are subsequently improved using a k-Nearest Neighbors model trained on spectroscopically confirmed quasars. The final catalog contains 1,295,502 quasar candidates with Gaia G band magnitudes brighter than 20.5.
@@
2.3 Pre-processing
We retrieve all light curves within the ZTF DR23 dataset using the bulk download option described in Section 12.c of the ZTF DR23 document [^3] https://irsa.ipac.caltech.edu/data/ZTF/docs/releases/dr23/ztf_release_notes_dr23.pdf. Each light curve contains the Heliocentric-based Modified Julian Date (HMJD), the point spread function (PSF) magnitude, 1-$\sigma$ photometric uncertainty, and corresponding photometric/image quality flags (catflags) for every observation. ZTF assigns unique object identifiers to sources based on combinations of right ascension (RA), declination (DEC), filter, field, and charge-coupled device (CCD) quadrant. We focus on the g and r bands data for their regular cadence and sky coverage (bellmZwickyTransientFacility2019), excluding the i band due to its insufficient sampling and limited source coverage.
Similar to grahamLightDarkSearching2023, we cross-match AGN positions from the combined catalog with ZTF DR23 sources using a $3^{\prime\prime}$ matching radius, resulting in approximately 3.6 million AGN light curves. To ensure sufficient sampling and data quality for flare detection, we require each light curve to contain more than 30 observations per band after removing epochs with catflags >0>0. This threshold is conservatively chosen given ZTF’s six-year baseline and is intended to provide sufficient temporal coverage for distinguishing flares from the AGN intrinsic variability. Figure 3 presents the distribution of valid ZTF detection counts per AGN in the g and r bands. Most AGNs have several hundred observations per band, ensuring robust coverage for variability analysis. The number of sources in each AGN catalog is shown in Table 1.
@@ 2.3 Pre-processing
3 method
3.1 Gaussian Processes
@@
where $\rho$ (variability amplitude) and $\tau$ (characteristic timescale) are hyperparameters. This kernel has proven to be effective in modeling AGN light curves (griffithsModelingMultiwavelengthVariability2021) and it is computationally friendly.
@@
For a given AGN light curve, the hyperparameters $\{\rho,\tau\}$ in the GP model can be optimized via maximum likelihood estimation. In mclaughlinUsingGaussianProcesses2024, GPs are employed to detect AGN flares by comparing the hyperparameters of individual AGNs to the overall hyperparameter distribution of the AGN population, with deviations potentially indicating the presence of a flare. However, since short-duration flares exert minimal influence on the parameter fits, this method is primarily sensitive to long-duration flares. In this work, we extend the approach from grahamLightDarkSearching2023 to detect flares of all durations across a large AGN sample.
@@
3.2 Initial Detection of AGN Flares
For each AGN with available ZTF data, we first fit its light curve (in flux space) using a Bayesian block representation (scargleSTUDIESASTRONOMICALTIME2013). This technique optimally segments the data into a series of discontinuous, piecewise constant components, facilitating the detection of significant local variations. To identify potential flares, we apply the hill-climbing procedure proposed by meyerCharacterizingGammaRayVariability2019. In this approach, peaks are identified as blocks that are higher than both the preceding and succeeding blocks, and these peaks are extended in both directions as long as the subsequent blocks remain lower. We define a flare interval in such a set of blocks as the segment where the flux exceeds 0.2 times the difference between the peak flux and the median flux of the entire light curve. If a flare interval consists of only one data point, we exclude it from further analysis.
@@ 3.2 Initial Detection of AGN Flares
Figure 5 illustrates the processing flow for a real AGN light curve, using the source J124942.29+344929.0 (RA=192.4262, Dec=34.8247) as an example. This AGN has a potential flare with $p_{\mathrm{flare}}=1.0$ and is ultimately retained in our final high-confidence flare sample. Gaussian Process modeling and related computations are performed using the Python package celerite, which is optimized for fast and scalable GP analysis on one-dimensional data (foreman-mackeyFastScalableGaussian2017). Following this procedure, we identify 28,504 AGN flare events, which are summarized in the AGN Flare Coarse Catalog - Version 1.0 (AGNFCC-V1.0).
@@
3.3 Refinement and Selection of High-Confidence Flares
While the initial flares identified using Gaussian Process modeling provide a valuable starting point, this coarse selection may still include false positives arising from irregular sampling or photometric noise. Moreover, given the large number of AGNs analyzed, some statistically significant events are expected to arised purely by chance, which is a manifestation of the look-elsewhere effect (vaughanFalsePeriodicitiesQuasar2016). To enhance the reliability of the flare sample, we apply a series of additional selection criteria designed to filter out ambiguous or low-significance events.

### v3

# A Systematic Search for Active Galactic Nucleus Flares in ZTF Data Release 23

@@ Abstract
software: celerite (Foreman-Mackey et al., 2017), astropy (Astropy Collaboration et al., 2013, 2018, 2022), matplotlib (Hunter, 2007), numpy (Harris et al., 2020), seaborn (Waskom, 2021), pandas (McKinney, 2010)
@@ I Introduction
Active galactic nuclei (AGNs) are extremely luminous sources powered by accretion onto supermassive black holes (SMBHs), and they exhibit significant variability across different wavelengths (Ulrich et al., 1997; Kelly et al., 2009). The optical light curves of AGNs often show stochastic brightness fluctuations on timescales from days to years, commonly modeled by a one-dimensional damped random walk (MacLeod et al., 2010; Kozłowski, 2016). These fluctuations can be attributed to various processes occurring near the supermassive black hole, such as accretion disk instabilities (Trèvese and Vagnetti, 2002), disk inhomogeneities (Dexter and Agol, 2010; Cai et al., 2016, 2018), or X-ray reprocessing (Kubota and Done, 2018), but the primary origin is still under debate. In addition to the general optical variability of AGN over different timescales, a subset of these systems exhibit transient and extreme events, known as AGN flares (Graham et al., 2017; Lawrence et al., 2016). These events represent a significant departure from the typical stochastic variability of AGN, yet their origins remain largely uncertain.
@@
A growing body of theoretical work suggests that many of these extreme flaring events may originate from stellar and compact objects embedded in AGN accretion disks, broadly referred to as accretion-modified stars (AMSs, Wang et al., 2021a, b; Liu et al., 2024). Within such dense and gas-rich environments, stars may evolve rapidly and end their lives as supernovae (Li et al., 2023), or be tidally disrupted by the central SMBH (Ryu et al., 2020) or stellar-mass black holes (Yang et al., 2022). These processes contribute to the metal enrichment of the AGN disk (Wang et al., 2023a) and the formation of compact objects. A large population of compact objects can form binaries (Rowan et al., 2023; Wang et al., 2024) and undergo mergers (McKernan et al., 2020a). The resulting coalescences can produce gravitational waves and potentially be accompanied by observable electromagnetic signals, including transients from white dwarf collisions (Zhang et al., 2023), kilonova emission from neutron star-neutron star or neutron star-black hole mergers (Ren et al., 2022), and flares from binary black hole mergers (Graham et al., 2020, 2023)
@@
These scenarios offer a unified framework for interpreting diverse AGN flaring phenomena and highlight the importance of searching for AGN flares as a means to probe the AMS populations. In addition, AGN flares serve as valuable diagnostics of accretion physics. Their timescales and amplitudes may constrain the properties of AGN disks and act as important probes for accretion rates (Ruan et al., 2016; Graham et al., 2017). Multi-epoch spectroscopy during the period of the flares further enables reverberation mapping to get the location of the emission region, helping constrain the structure of the AGN disk (Payne et al., 2023; Cabrera et al., 2024).
@@
Recognizing their scientific potential, several systematic searches have been conducted to identify AGN flares (Lawrence et al., 2016; Graham et al., 2017), utilizing data from large surveys like Panoramic Survey Telescope and Rapid Response System (Pan-STARRS, Kaiser et al., 2010), Sloan Digital Sky Survey (SDSS, Sesar et al., 2007), and Catalina Real-time Transient Survey (CRTS, Drake et al., 2009). These studies have primarily focused on detecting long-duration, high-amplitude variability events, primarily due to the limited observational cadence of these surveys.
@@
The emergence of high-cadence time-domain facilities, such as the Zwicky Transient Facility (ZTF, Bellm et al., 2019; Graham et al., 2019), the Wide Field Survey Telescope (WFST, Wang et al., 2023b), and the Vera C. Rubin observatory (Collaboration et al., 2017; Ivezić et al., 2019), would be able to capture much more AGN flares across a wide range of timescales. These advances would facilitate the discovery and characterization of AGN flares, offering deeper insights into the physical mechanisms driving such outbursts.
@@
To fully utilize the data from modern high-cadence surveys, recent studies have proposed the use of Gaussian Processes (GPs) for the detection of AGN flares (McLaughlin et al., 2024). Graham et al. (2023) applied a GP-based technique to separate flaring events from typical AGN variability, while He et al. (2025) further demonstrated that applying such methods to larger datasets significantly improves the reliability of flare identification. Motivated by these developments, we carry out a systematic search for AGN flares in ZTF Data Release 23 (DR23) using a GP-based approach, aiming to construct the most comprehensive catalog of such events to date.
@@
The outline of this paper is as follows. In Section II, we describe the ZTF dataset and the AGN catalogs used in this work. In Section III, we introduce the flare detection method, including Gaussian Process modeling and the selection criteria used to construct two catalogs: the AGN Flare Coarse Catalog (AGNFCC) and the AGN Flare Refined Catalog (AGNFRC). In Section IV, we present the statistical properties of the detected flares. In Section V, we discuss their potential origins. Finally, we summarize our main findings in Section VI.
II data
II.1 ZTF DR23
@@
The Zwicky Transient Facility (ZTF) is a wide-field time-domain survey utilizing a 47 deg2 field-of-view camera mounted on the Palomar 48-inch Samuel Oschin Schmidt telescope. Designed to systematically explore transient and variable phenomena with high cadence and sensitivity (Bellm et al., 2018; Graham et al., 2019), ZTF has been operating since March 2018, covering approximately $3\pi$ sr across the Northern sky (Bellm et al., 2019). For the Northern sky of declination $31^{\circ}$, ZTF operates two primary observing programs: the Survey, which targets fields with $|b|<7^{\circ}$ at a 1-day cadence, and the Northern Sky Survey, which observes fields with $|b|>7^{\circ}$ at a 3-day cadence. Both surveys acquire data in the g and r bands, providing dual-band light curves for each source.
@@ II.2 AGN Catalogs
Million Quasars Catalog (Milliquas, v8, Flesch, 2023): A comprehensive catalog of spectroscopically confirmed AGNs and high-confidence AGN candidates ($p_{\mathrm{QSO}}\geq 99\%$), compiled from all published papers up to 2023 June 30. It contains 1,021,800 sources, with $90\%$ having spectroscopic redshifts and the remainder utilizing photometric redshifts. This catalog incorporates astrometric corrections to historical positional errors and integrates high-precision multi-survey data, delivering the most precise locations.
DESI AGN Catalog (Collaboration et al., 2025): The Dark Energy Spectroscopic Instrument (DESI) has conducted the most extensive extragalactic spectroscopic redshift survey to date. Its first Data Release (DR1) provides high-confidence redshifts for approximately 18 million objects across over 9,000 deg2. In this work, we use the subset of objects labeled with SPECTYPE = ‘QSO’ in the DESI value-added catalog [^1] https://data.desi.lbl.gov/public/dr1/vac/dr1/agnqso (Juneau et al 2025, in preparation), which includes a total of 1,772,694 AGNs. This represents the largest spectroscopic AGN sample available to date, making it a valuable source for statistical studies of AGN.
LAMOST AGN Catalog (Ai et al., 2016; Dong et al., 2018; Yao et al., 2019; Jin et al., 2023): A set of spectroscopically confirmed AGNs from the Large Sky Area Multi-Object Fiber Spectroscopic Telescope (LAMOST) quasar survey conducted between 2011 and 2021. The full sample contains 55,636 quasars, of which over 30,000 are independently discovered. All spectra are recalibrated using SDSS or Pan-STARRS photometry, and the key spectral properties, including emission line widths, continuum luminosities, and black hole masses, are homogeneously measured.
Quaia AGN Catalog (Storey-Fisher et al., 2024): An all-sky catalog constructed from sources identified as quasar candidates in Gaia Data Release 3 (Bailer-Jones et al., 2023), initially comprising 6,649,162 candidates with redshift estimates derived from the low-resolution blue photometer/red photometer spectra. To improve the purity of the sample, these candidates are cross-matched with the unWISE infrared catalog (Schlafly et al., 2019) and refined by applying cuts based on color and proper motion to reduce contamination. Redshift estimates are subsequently improved using a k-Nearest Neighbors model trained on spectroscopically confirmed quasars. The final catalog contains 1,295,502 quasar candidates with Gaia G band magnitudes brighter than 20.5.
@@
II.3 Pre-processing
We retrieve all light curves within the ZTF DR23 dataset using the bulk download option described in Section 12.c of the ZTF DR23 document [^3] https://irsa.ipac.caltech.edu/data/ZTF/docs/releases/dr23/ztf_release_notes_dr23.pdf. Each light curve contains the Heliocentric-based Modified Julian Date (HMJD), the point spread function (PSF) magnitude, 1-$\sigma$ photometric uncertainty, and corresponding photometric/image quality flags (catflags) for every observation. ZTF assigns unique object identifiers to sources based on combinations of right ascension (RA), declination (DEC), filter, field, and charge-coupled device (CCD) quadrant. We focus on the g and r bands data for their regular cadence and sky coverage (Bellm et al., 2019), excluding the i band due to its insufficient sampling and limited source coverage.
Similar to Graham et al. (2023), we cross-match AGN positions from the combined catalog with ZTF DR23 sources using a $3^{\prime\prime}$ matching radius, resulting in approximately 3.6 million AGN light curves. To ensure sufficient sampling and data quality for flare detection, we require each light curve to contain more than 30 observations per band after removing epochs with catflags >0>0. This threshold is conservatively chosen given ZTF’s six-year baseline and is intended to provide sufficient temporal coverage for distinguishing flares from the AGN intrinsic variability. Figure 3 presents the distribution of valid ZTF detection counts per AGN in the g and r bands. Most AGNs have several hundred observations per band, ensuring robust coverage for variability analysis. The number of sources in each AGN catalog is shown in Table 1.
@@ II.3 Pre-processing
III method
III.1 Gaussian Processes
@@
where $\rho$ (variability amplitude) and $\tau$ (characteristic timescale) are hyperparameters. This kernel has proven to be effective in modeling AGN light curves (Griffiths et al., 2021) and it is computationally friendly.
@@
For a given AGN light curve, the hyperparameters $\{\rho,\tau\}$ in the GP model can be optimized via maximum likelihood estimation. In McLaughlin et al. (2024), GPs are employed to detect AGN flares by comparing the hyperparameters of individual AGNs to the overall hyperparameter distribution of the AGN population, with deviations potentially indicating the presence of a flare. However, since short-duration flares exert minimal influence on the parameter fits, this method is primarily sensitive to long-duration flares. In this work, we extend the approach from Graham et al. (2023) to detect flares of all durations across a large AGN sample.
@@
III.2 Initial Detection of AGN Flares
For each AGN with available ZTF data, we first fit its light curve (in flux space) using a Bayesian block representation (Scargle et al., 2013). This technique optimally segments the data into a series of discontinuous, piecewise constant components, facilitating the detection of significant local variations. To identify potential flares, we apply the hill-climbing procedure proposed by Meyer et al. (2019). In this approach, peaks are identified as blocks that are higher than both the preceding and succeeding blocks, and these peaks are extended in both directions as long as the subsequent blocks remain lower. We define a flare interval in such a set of blocks as the segment where the flux exceeds 0.2 times the difference between the peak flux and the median flux of the entire light curve. If a flare interval consists of only one data point, we exclude it from further analysis.
@@ III.2 Initial Detection of AGN Flares
Figure 5 illustrates the processing flow for a real AGN light curve, using the source J124942.29+344929.0 (RA=192.4262, Dec=34.8247) as an example. This AGN has a potential flare with $p_{\mathrm{flare}}=1.0$ and is ultimately retained in our final high-confidence flare sample. Gaussian Process modeling and related computations are performed using the Python package celerite, which is optimized for fast and scalable GP analysis on one-dimensional data (Foreman-Mackey et al., 2017). Following this procedure, we identify 28,504 AGN flare events, which are summarized in the AGN Flare Coarse Catalog - Version 1.0 (AGNFCC-V1.0).
@@
III.3 Refinement and Selection of High-Confidence Flares
While the initial flares identified using Gaussian Process modeling provide a valuable starting point, this coarse selection may still include false positives arising from irregular sampling or photometric noise. Moreover, given the large number of AGNs analyzed, some statistically significant events are expected to arised purely by chance, which is a manifestation of the look-elsewhere effect (Vaughan et al., 2016). To enhance the reliability of the flare sample, we apply a series of additional selection criteria designed to filter out ambiguous or low-significance events.